OpenStudy (anonymous):

What is the quotient: (2x3 – 13x2 + 16x – 5) ÷ (x – 5)

OpenStudy (anonymous):

By long division, quotient is 2x^2-3x+1

OpenStudy (anonymous):

(x-5)(x-1)(x-0.5) / (X-5) = (X-1)((x-0.5)

OpenStudy (anonymous):

(2x^3-13x2+16x-5) = (x-5)(x-1)(x-0.5)